A stone and mass m is thrown from earth`s surface at an angle theta to the horizontal with an initial velocity u. Find out the average force just t second after the beginning of motion

Grade:12th pass

1 Answers

Vikas TU
7 years ago
After t seconds, final velocites in x and y direcns respectively are:
Vx = ucostheta
Vy = usintheta – gt
Net velocity = root((ucosthetha)^2 + (usintheta – gt)^2)
dV/dt = (usintheta – gt)*(-g)/((ucosthetha)^2 + (usintheta – gt)^2)
 
Force = m*a = > m*dV/dt = > m*(usintheta – gt)*(-g)/((ucosthetha)^2 + (usintheta – gt)^2) Netwon.